OpenStudy (anonymous):

Solving absolute equations. |m+5|+1.5=2

OpenStudy (anonymous):

m+5+1.5=2 m+6.5=2 m=2-6.5 m=-4.5 -m-5+1.5=2 -m-3.5=2 -3.5-2=m -5.5=m